OpenStudy (anonymous):

The measures of 5 of the interior angles of a hexagon are: 130°, 120°, 80°, 160°, and 155°. What is the measure of the largest exterior angle? 100° 105° 95° 90°

jimthompson5910 (jim_thompson5910):

hint: all the interior angles must add to 720 degrees how am I getting 720? by using the formula S = 180(n-2) and plugging in n = 6 S = 180(n-2) S = 180(6-2) S = 180(4) S = 720

jimthompson5910 (jim_thompson5910):

Once you know all the interior angles, you can find the exterior angles of each interior angle.

OpenStudy (anonymous):

yes I know that part but idk the exterior part

jimthompson5910 (jim_thompson5910):

if x is any interior angle and y is its corresponding exterior angle, then x+y = 180

jimthompson5910 (jim_thompson5910):

so for instance, the first interior angle is 130 degrees, so x = 130 x+y = 180 130+y = 180 y = 180 - 130 y = 50 so the exterior angle is 50 degrees

OpenStudy (anonymous):

thank you that really helped me:)

jimthompson5910 (jim_thompson5910):

you're welcome